if the two sidea of a rectangle change so the new area is double the original area,and one side is increased by 50%,what percentage increase or decrease of the other side????
Answer by HyperBrain(694)

Let A=area of the original rectangle
x=length of the orig. rectangle
y=width of the orig. rectangle
Then
A=x*y
If the new area is double the orig. rectangle
Then
2*x*y=area of the new rectangle
if you increase one side (let's say the length) by 50%, you are adding x/2
so the new length is x+(x/2)=3x/2
Let W=new width
Then
2*x*y=(3x/2)W
4*x*y=3xW multiply by 2 both sides
4y=3W divide both sides by x
W=4y/3=y +(y/3)
1/3 is 33 1/3%.
Therefore, the width is increased by 33 1/3%.
Power up,
HyperBrain!
Help us promote our advocacy no straws, no plastics, no styros.
Visit these sites:
http://uselessplastic.wordpress.com
http://uselessplastic.multiply.com